OpenStudy (anonymous):

Mario answered 75% of the 40 problems on his history homework correctly. How many problems on his homework did he answer correctly 30 19 28 10

OpenStudy (anonymous):

40(.75) or 40*(3/4)= 120/4 = 30 He answered correctly 30 questions

OpenStudy (anonymous):

.75*40

OpenStudy (anonymous):

OH come on, seriously? 40*0.75=30.

OpenStudy (anonymous):

dont be rude -___-

OpenStudy (anonymous):

Im reviewing over past grades and I forgot most Of it so dont be rude bro....